# SAMix: Calibrated and Accurate Continual Learning via Sphere-Adaptive Mixup and Neural Collapse

### v1

@@ 3.3 Sphere-Adaptive Mixup
Our new mixup method utilizes mixed samples and prototypes to improve smooth boundaries in representation learning and enhance calibration. Each batch is drawn from the current task data $\mathcal{D}_{t}$ and the memory buffer $\mathcal{M}$. Given two random samples $(\mathbf{x}_{i},y_{i})$ and $(\mathbf{x}_{j},y_{j})$ with corresponding prototypes $\mathbf{p}_{y_{i}}$ and $\mathbf{p}_{y_{j}}$, we generate a mixed sample $\tilde{\mathbf{x}}_{ij}$ as in Eq. (1). Unlike original mixup (Zhang et al. 2018), which combines target labels as in Eq. (2), we generate new prototypes based on the mixing coefficient $\lambda$. To generate mixed prototypes, several works (Yang et al. 2022; Fisher, Meng, and Papyan 2024) use linear interpolation, defined as:
@@
Figure 2: Linear interpolation and Slerp.
@@
$=\lambda^{2}\|\mathbf{p}_{y_{i}}\|_{2}^{2}+2\lambda(1-\lambda)\langle\mathbf{p}_{y_{i}}\cdot\mathbf{p}_{y_{j}}\rangle$
$\qquad+{(1-\lambda)^{2}}{\|\mathbf{p}_{y_{j}}\|_{2}^{2}}$ (4)
@@
To ensure that mixed prototypes always remain on the unit hypersphere, we use Spherical Linear Interpolation (Slerp) (Shoemake 1985) instead of linear interpolation. A geometric visualization of this approach is shown in Figure 2b. The hypersphere-interpolated prototype $\tilde{\mathbf{p}}_{ij}$ is:
@@ 3.4 Integration of SAMix into NC-based Methods
We note that mixed samples from SAMix are used only in $\mathcal{L}_{plas}$ and not in $\mathcal{L}_{stab}$. On the first hand, mixed samples are designed to enhance plasticity for new tasks (smoother transitions between tasks). On the other hand, $\mathcal{L}_{stab}$ focuses on preserving old knowledge: when using relation-based distillation techniques, such as instance-wise or instance-prototype relationships, only the fixed prototypes and original samples should be used, not the mixed ones. This approach aligns with the work in (Beyer et al. 2022), which suggests that a “patient” teacher is ideal for knowledge distillation, meaning mixed samples, and thus mixed class prototypes are unnecessary for distillation.
SAMix is designed to complement any NC-based method that relies on fixed prototypes. In this work, we investigate SAMix with two prominent NC-based plastic losses: dot-regression (DR) (Yang et al. 2022) and the recent focal neural collapse contrastive (FNC2) (Dang et al. 2025) which are respectively defined in Eq. (7) and Eq. (8):
@@
$\mathcal{L}_{FNC^{2}}=-\sum_{i=1}^{2N}\frac{1}{|P(i)|+1}\Bigg($ $\sum_{\mathbf{z}_{j}\in P(i)}(1-c_{ij})^{\gamma}\log(c_{ij})$
$+(1-r_{i})^{\gamma}\log(r_{i})\Bigg)$ (8)
@@
Training batch composition. Each batch includes $2N$ normal samples (see Sec. 3.1) and $2N$ mixed samples ($4N$ total). Despite the increased batch size, our method achieves significant performance gains with little extra time and remains effective even with smaller batches (see Appendix).
@@
Stability loss. For $\mathcal{L}_{stab}$, we use hardness-softness distillation (HSD) (Dang et al. 2025), which surpasses instance-wise relation distillation (Cha, Lee, and Shin 2021) and sample-prototype relation distillation (S-PRD) (Dang et al. 2025) in preserving old knowledge (see Appendix for loss definitions). We refer to the approach in (Dang et al. 2025) as Focal Contrastive Neural Collapse Continual Learning (FC-NCCL). Besides, we propose a novel method using DR for plasticity and HSD for stability, which we refer to Tightly Aligned Neural Collapse Continual Learning (TA-NCCL).
@@ 3.5 Expected Outcome Hypothesis
(a) H1: Performance without SAMix. FC-NCCL is expected to outperform TA-NCCL, as it better preserves within-class distributions during representation learning (Dang et al. 2025). It is also anticipated to outperform TA-NCCL in calibration, since its FNC2 loss already integrates temperature scaling and focal loss, two well-established calibration techniques (Guo et al. 2017; Lin et al. 2017).
(b) H2: Limited benefit of SAMix for FNC2. We hypothesize that SAMix provides limited benefits for FNC2 in FC-NCCL when FNC2 is applied to normal samples and DR to mixed ones. As a contrastive loss, FNC2 benefits from more negative samples (Chen et al. 2020), which SAMix does not provide. Therefore, we predict that any performance gains from integrating SAMix primarily come from the DR loss, as it better leverages mixed samples. In contrast, FNC2 relies on a large number of negative samples, which SAMix does not address.
(c) H3: Significant benefit of SAMix for DR. In contrast, we expect DR, which operates on mixed samples, to fully benefit from SAMix. Unlike FNC2, which considers both sample-sample and sample-prototype contrastive relationships for flexible clustering, DR only focuses on sample-prototype relations, resulting in distinct but less flexible clusters. However, when SAMix is applied, the decision boundaries become smoother. As mentioned in (Fisher, Meng, and Papyan 2024), using mixup can create new configurations with mixed samples, significantly enhancing representation learning. We expect this will allow DR to learn more robust representations, improve generalization, and perform better on both the current task and across tasks in CL, while also mitigating forgetting.
@@ 4 Experiments
Buffer Dataset Venue Seq-Cifar-100 Seq-Tiny-ImageNet Seq-Cifar-10
@@
Table 1: Comparison of our method with supervised baselines at memory sizes 200 and 500, averaged over 5 trials from 5 different seeds (best results in bold). $\blacktriangle$ indicates performance improvement when applying SAMix. (Ours: TA-NCCL method).
Figure 3: Results on Seq-Cifar100 with the Class-IL setting.
@@ 4.1 Datasets and Implementation details
Datasets. We use three standard CL datasets widely used in literature for fair comparison under Class-IL and Task-IL settings (3.1): Seq-Cifar-10 and Seq-Cifar-100 (Krizhevsky and Hinton 2009), and Seq-Tiny-ImageNet (Le and Yang 2015). Seq-Cifar-10 has 5 tasks with 2 classes each; Seq-Cifar-100 and Seq-Tiny-ImageNet each have 5 and 10 tasks with 20 classes per task, respectively.
@@
Implementation details. We utilize ResNet-18 (He et al. 2016) as the backbone, omitting the final layer as in (Cha, Lee, and Shin 2021; Fini et al. 2022; Dang et al. 2025; Li et al. 2024a). Following (Chen et al. 2020; Zbontar et al. 2021; Cha, Lee, and Shin 2021), we add a two-layer projection MLP $g^{t}$ to map backbone features into a $d$-dimensional space ($d=128$ for Seq-Cifar-10/100, $d=256$ for Seq-Tiny-ImageNet). For distillation, as in (Fini et al. 2022), we use a predictor $h^{t}$ with the same architecture as $g^{t}$ to align current and past feature spaces. Besides, we use buffer sizes of 0, 200, and 500. For evaluation, following (Cha, Lee, and Shin 2021; Wen et al. 2024; Dang et al. 2025), we train a classifier $s^{t}$ on top of the frozen encoder $f^{t}$ for 100 epochs, using samples from the last-task training dataset $\mathcal{D}_{t}$ and buffer $\mathcal{M}$. As in (Cha, Lee, and Shin 2021), no buffer is used during representation learning in the memory-free setting, but 200 auxiliary samples are still needed to only train the classifier for all datasets. Experiments are conducted on a NVIDIA Quadro GV100 GPU with 256GB RAM, Ubuntu 20.04, and PyTorch 1.12.1 (Paszke et al. 2019).
@@
Considered methods. We evaluate SAMix across the SOTA NC-based methods: FC-NCCL (Dang et al. 2025) and TA-NCCL. We also consider recent baselines, including CCLIS (Li et al. 2024a), CILA (Wen et al. 2024), and other well-known supervised methods such as ER (Riemer et al. 2019), DER (Buzzega et al. 2020), Co2L (Cha, Lee, and Shin 2021), and GCR (Tiwari et al. 2022).
@@
Evaluation metrics. As in prior work (Cha, Lee, and Shin 2021; Fini et al. 2022; Wen et al. 2024; Dang et al. 2025; Li et al. 2024a), we report Average Accuracy (AA) (Chaudhry et al. 2018) across all tasks on the test dataset after learning the final task $T$: $\text{AA}=\frac{1}{T}\sum_{k=1}^{T}\text{A}_{T,k}$. In addition, we use Average Forgetting (Chaudhry et al. 2018) to measure how much the model forgets previous tasks (see Appendix).
@@
Calibration metrics. To assess network calibration, we use the metrics from (Guo et al. 2017). Predictions are divided into $M$ interval equal-sized bins, with $B_{m}$ representing the set of samples whose prediction scores (the winning softmax score) fall into bin $m$. The accuracy and confidence of $B_{m}$ are defined as: $\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})$; $\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$, where $\hat{p}_{i}$ is the confidence (winning score) of sample $i$. The Expected Calibration Error (ECE) is defined as: $\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}|\text{acc}(B_{m})-\text{conf}(B_{m})|$. We also use the Overconfidence Error (OE) metric (Thulasidasan et al. 2019), essential for high-risk applications where confidently incorrect predictions are critical: $\text{OE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}[\text{conf}(B_{m})\times\max(\text{conf}(B_{m})-\text{acc}(B_{m}),0)]$. This metric heavily penalizes overconfident predictions when confidence exceeds accuracy.
@@ 4.2 Main Results
However, on Seq-Cifar-100, the trend reverses, with DR outperforming FNC2 entirely, which contradicts the initial performance hypothesis. We attribute this to Cifar-100’s higher inter-class similarity (Bertinetto et al. 2019), stemming from its 100 classes grouped into 20 semantically related superclasses (e.g., vehicles, animals). In contrast, Tiny ImageNet contains 200 more diverse and ungrouped classes. This similarity in Cifar-100 may lead to greater class confusion, limiting the advantages of FC-NCCL. Moreover, as shown in Figure 3a, FNC2 achieves lower ECE5,5 on the current task, indicating better calibration than DR. However, with buffer usage, FC-NCCL shows higher OE5,5 than TA-NCCL (see Figure 3b).
b) H2: Limited benefit of SAMix for FNC2. This hypothesis holds partially, as SAMix yields slight improvements on Seq-Tiny-ImageNet and even degrades performance on Seq-Cifar-10 with buffer sizes of 0 and 200. In methods using plasticity loss FNC2, applying DR to mixed samples often shift the model’s focus toward blended artifacts rather than meaningful features, leading to forgetting. In contrast, on Seq-Cifar-100, SAMix significantly enhances FC-NCCL performance, which is attributed to the dataset’s high inter-class similarity (Bertinetto et al. 2019). In this case, SAMix refines class boundaries and boosts performance. Without it, performance drops substantially compared to DR-based, which further highlights its benefits.
@@ 4.3 Ablation Studies
Table 2: Comparison between Slerp and linear interpolation on Seq-Cifar-100. $\blacktriangle$ indicates that Slerp performs better.
Effectiveness of using slerp in SAMix. We report ablation results in Table 2, comparing Slerp and linear interpolation when applying SAMix to both FC-NCCL and TA-NCCL on Seq-Cifar-100 with buffer sizes of 0, 200, and 500. Slerp yields better calibration and lower AOE in all cases than linear interpolation. Moreover, it also consistently improves accuracy at larger distances. These results support the use of Slerp in SAMix and align with our analysis in Sec. 3.3.

### v2

@@ 3.3 Sphere-Adaptive Mixup
Our new mixup method utilizes mixed samples and prototypes to improve smooth boundaries in representation learning and enhance calibration. Each batch is drawn from the current task data $\mathcal{D}_{t}$ and the memory buffer $\mathcal{M}$. Given two random samples $(\mathbf{x}_{i},y_{i})$ and $(\mathbf{x}_{j},y_{j})$ with corresponding prototypes $\mathbf{p}_{y_{i}}$ and $\mathbf{p}_{y_{j}}$, we generate a mixed sample $\tilde{\mathbf{x}}_{ij}$ as in Eq. (1). Unlike original mixup [mixup_2018], which combines target labels as in Eq. (2), we generate new prototypes based on the mixing coefficient $\lambda$. To generate mixed prototypes, several works [yang_induce_nc_2022, fisher2024pushingboundariesmixupsinfluence] use linear interpolation, defined as:
@@
Figure 2: Linear interpolation and Slerp
@@
$=\lambda^{2}\|\mathbf{p}_{y_{i}}\|_{2}^{2}+2\lambda(1-\lambda)\langle\mathbf{p}_{y_{i}}\cdot\mathbf{p}_{y_{j}}\rangle+{(1-\lambda)^{2}}{\|\mathbf{p}_{y_{j}}\|_{2}^{2}}$ (6)
@@
To ensure that mixed prototypes always remain on the unit hypersphere, we use Spherical Linear Interpolation (Slerp) [shoemake_slerp_1985] instead of linear interpolation. A geometric visualization of this approach is shown in Figure 2b. The hypersphere-interpolated prototype $\tilde{\mathbf{p}}_{ij}$ is:
@@ 3.4 Integration of SAMix into NC-based Methods
We note that mixed samples from SAMix are used only in $\mathcal{L}_{plas}$ and not in $\mathcal{L}_{stab}$. On the first hand, mixed samples are designed to enhance plasticity for new tasks (smoother transitions between tasks). On the other hand, $\mathcal{L}_{stab}$ focuses on preserving old knowledge: when using relation-based distillation techniques, such as instance-wise or instance-prototype relationships, only the fixed prototypes and original samples should be used, not the mixed ones. This approach aligns with the work in [beyer2022knowledgedistillationgoodteacher], which suggests that a “patient” teacher is ideal for knowledge distillation, meaning mixed samples, and thus mixed class prototypes are unnecessary for distillation. Moreover, we empirically experimented with incorporating SAMix into the stability loss, but observed no performance improvement while incurring a noticeable increase in training time. Therefore, we exclude SAMix from $\mathcal{L}_{stab}$, resulting in a favorable trade-off between additional training time and performance gains, as discussed in Sec. 4.3.2.
SAMix is designed to complement any NC-based method that relies on fixed prototypes. In this work, we investigate SAMix with two prominent NC-based plastic losses: dot-regression (DR) [yang_induce_nc_2022] and the recent focal neural collapse contrastive (FNC2) [anh_wacv_2025] which are respectively defined in Eq. (9) and Eq. (10):
@@
$\mathcal{L}_{FNC^{2}}=-\sum_{i=1}^{2N}\frac{1}{|P(i)|+1}\Bigg($ $\sum_{\mathbf{z}_{j}\in P(i)}(1-c_{ij})^{\gamma}\log(c_{ij})+(1-r_{i})^{\gamma}\log(r_{i})\Bigg)$ (10)
@@
Training batch composition. Each batch includes $2N$ normal samples (see Sec. 3.1) and $2N$ mixed samples ($4N$ total). Despite the increased batch size, our method achieves significant performance gains with little extra time (see Sec. 4.3.2) and remains effective even with smaller batches (see Sec. 4.3.3).
@@
Stability loss. For $\mathcal{L}_{stab}$, we use hardness-softness distillation (HSD) [anh_wacv_2025], which surpasses instance-wise relation distillation [co2l] and sample-prototype relation distillation (S-PRD) [anh_wacv_2025] in preserving old knowledge (see Appendix B for loss definitions). We refer to the approach in [anh_wacv_2025] as Focal Contrastive Neural Collapse Continual Learning (FC-NCCL). Besides, we propose a novel method using DR for plasticity and HSD for stability, which we refer to Tightly Aligned Neural Collapse Continual Learning (TA-NCCL).
@@ 3.5 Expected Outcome Hypothesis
(a) H1: Performance without SAMix. FC-NCCL is expected to outperform TA-NCCL, as it better preserves within-class distributions during representation learning [anh_wacv_2025]. It is also anticipated to outperform TA-NCCL in calibration, since its FNC2 loss already integrates temperature scaling and focal loss, two well-established calibration techniques [guo2017_calibration_nn_2017, lin2018focallossdenseobject].
(b) H2: Limited benefit of SAMix for FNC2. We hypothesize that SAMix provides limited benefits for FNC2 in FC-NCCL when FNC2 is applied to normal samples and DR to mixed ones. As a contrastive loss, FNC2 benefits from more negative samples [simclr-chen20j], which SAMix does not provide. Therefore, we predict that any performance gains from integrating SAMix primarily come from the DR loss, as it better leverages mixed samples. In contrast, FNC2 relies on a large number of negative samples, which SAMix does not address.
(c) H3: Significant benefit of SAMix for DR. In contrast, we expect DR, which operates on mixed samples, to fully benefit from SAMix. Unlike FNC2, which considers both sample-sample and sample-prototype contrastive relationships for flexible clustering, DR only focuses on sample-prototype relations, resulting in distinct but less flexible clusters. However, when SAMix is applied, the decision boundaries become smoother. As mentioned in [fisher2024pushingboundariesmixupsinfluence], using mixup can create new configurations with mixed samples, significantly enhancing representation learning. We expect this will allow DR to learn more robust representations, improve generalization, and perform better on both the current task and across tasks in CL, while also mitigating forgetting.
@@ 4 Experiments
Table 1: Comparison of our method with supervised baselines at memory sizes 200 and 500, averaged over 5 trials from 5 different seeds (best results in bold). $\blacktriangle$ indicates performance improvement when applying SAMix. (Ours: TA-NCCL method)
Buffer Dataset Seq-Cifar-100 Seq-Tiny-ImageNet Seq-Cifar-10
@@
Figure 3: Results on Seq-Cifar100 with the Class-IL setting
@@ 4.1 Datasets and Implementation details
Datasets. We use three standard CL datasets widely used in literature for fair comparison under Class-IL and Task-IL settings (3.1): Seq-Cifar-10 and Seq-Cifar-100 [krizhevsky2009learning], and Seq-Tiny-ImageNet [le2015tinyiv]. Seq-Cifar-10 has 5 tasks with 2 classes each; Seq-Cifar-100 and Seq-Tiny-ImageNet each have 5 and 10 tasks with 20 classes per task, respectively.
@@
Implementation details. We utilize ResNet-18 [resnet18] as the backbone, omitting the final layer as in [co2l, fini2022cassle, anh_wacv_2025, li_cl_important_sampling_2024]. Following [simclr-chen20j, zbontar2021barlowtwinsselfsupervisedlearning, co2l], we add a two-layer projection MLP $g^{t}$ to map backbone features into a $d$-dimensional space ($d=128$ for Seq-Cifar-10/100, $d=256$ for Seq-Tiny-ImageNet). For distillation, as in [fini2022cassle], we use a predictor $h^{t}$ with the same architecture as $g^{t}$ to align current and past feature spaces. Besides, we use buffer sizes of 0, 200, and 500. For evaluation, following [co2l, wen2024provablecontrastivecontinuallearning, anh_wacv_2025], we train a classifier $s^{t}$ on top of the frozen encoder $f^{t}$ for 100 epochs, using samples from the last-task training dataset $\mathcal{D}_{t}$ and buffer $\mathcal{M}$. As in [co2l], no buffer is used during representation learning in the memory-free setting, but 200 auxiliary samples are still needed to only train the classifier for all datasets. Experiments are conducted on a NVIDIA Quadro GV100 GPU with 256GB RAM, Ubuntu 20.04, and PyTorch 1.12.1 [paszke2019pytorchimperativestylehighperformance].
@@
Considered methods. We evaluate SAMix across the SOTA NC-based methods: FC-NCCL [anh_wacv_2025] and TA-NCCL. We also consider recent baselines, including CCLIS [li_cl_important_sampling_2024], CILA [wen2024provablecontrastivecontinuallearning], and other well-known supervised methods such as ER [riemer2019learninglearnforgettingmaximizing], DER [buzzega2020darkexperiencegeneralcontinual], Co2L [co2l], and GCR [tiwari2022gcrgradientcoresetbased].
@@
Evaluation metrics. As in prior work [co2l, fini2022cassle, wen2024provablecontrastivecontinuallearning, anh_wacv_2025, li_cl_important_sampling_2024], we report Average Accuracy (AA) [chaudhry_2018] across all tasks on the test dataset after learning the final task $T$: $\text{AA}=\frac{1}{T}\sum_{k=1}^{T}\text{A}_{T,k}$. In addition, we use Average Forgetting [chaudhry_2018] to measure how much the model forgets previous tasks (see Appendix D.1).
@@
Calibration metrics. To assess network calibration, we use the metrics from [guo2017_calibration_nn_2017]. Predictions are divided into $M$ interval equal-sized bins, with $B_{m}$ representing the set of samples whose prediction scores (the winning softmax score) fall into bin $m$. The accuracy and confidence of $B_{m}$ are defined as: $\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})$; $\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$, where $\hat{p}_{i}$ is the confidence (winning score) of sample $i$. The Expected Calibration Error (ECE) is defined as: $\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}|\text{acc}(B_{m})-\text{conf}(B_{m})|$. We also use the Overconfidence Error (OE) metric [thulasidasan_on_mixup_training_2020], essential for high-risk applications where confidently incorrect predictions are critical: $\text{OE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}[\text{conf}(B_{m})\times\max(\text{conf}(B_{m})-\text{acc}(B_{m}),0)]$. This metric heavily penalizes overconfident predictions when confidence exceeds accuracy.
@@ 4.2 Main Results
However, on Seq-Cifar-100, the trend reverses, with DR outperforming FNC2 entirely, which contradicts the initial performance hypothesis. We attribute this to Cifar-100’s higher inter-class similarity [bertinetto2018metalearning], stemming from its 100 classes grouped into 20 semantically related superclasses (e.g., vehicles, animals). In contrast, Tiny ImageNet contains 200 more diverse and ungrouped classes. This similarity in Cifar-100 may lead to greater class confusion, limiting the advantages of FC-NCCL. Moreover, as shown in Figure 3a, FNC2 achieves lower ECE5,5 on the current task, indicating better calibration than DR. However, with buffer usage, FC-NCCL shows higher OE5,5 than TA-NCCL (see Figure 3b).
b) H2: Limited benefit of SAMix for FNC2. This hypothesis holds partially, as SAMix yields slight improvements on Seq-Tiny-ImageNet and even degrades performance on Seq-Cifar-10 with buffer sizes of 0 and 200. The degradation on Seq-Cifar-10 is mainly due to its setting, where each task contains only two classes. In memory-free or low-memory scenarios, mixed samples thus lack diversity, as they primarily come from these two classes. Furthermore, the number of mixed samples per batch equals the batch size, making it challenging for FNC2, which relies on the “push and pull” mechanism, to achieve the alignment and uniformity properties [wang2022understandingcontrastiverepresentationlearning] required for effective contrastive representation learning. As a result, FNC2 struggles to pull same-class samples together and learn distinct, cluster-based representations.
Moreover, in methods using plasticity loss FNC2, applying DR to mixed samples often shift the model’s focus toward blended artifacts rather than meaningful features, leading to forgetting. In contrast, on Seq-Cifar-100, SAMix significantly enhances FC-NCCL performance, which is attributed to the dataset’s high inter-class similarity [bertinetto2018metalearning]. In this case, SAMix refines class boundaries and boosts performance. Without it, performance drops substantially compared to DR-based, which further highlights its benefits.
@@ 4.3 Ablation Studies
Table 2: Comparison between Slerp and linear interpolation on Seq-Cifar-100. $\blacktriangle$ indicates that Slerp performs better
@@
4.3.1 Effectiveness of using slerp in SAMix
We report ablation results in Table 2, comparing Slerp and linear interpolation when applying SAMix to both FC-NCCL and TA-NCCL on Seq-Cifar-100 with buffer sizes of 0, 200, and 500. Slerp yields better calibration and lower AOE in all cases than linear interpolation. Moreover, it also consistently improves accuracy at larger distances. These results support the use of Slerp in SAMix and align with our analysis in Sec. 3.3.
4.3.2 Extra time when using SAMix
Although our method increase batch size (see Sec. 3.4), mixed samples are used only in the plasticity loss (DR), not in distillation or contrastive losses like FNC2 to avoid disrupting the mixed-up configuration (see Appendix proof A). Thus, mixed samples are not used as regular inputs, keeping the impact on training time minimal. Table 3 shows the extra time when applying SAMix to FC-NCCL and TA-NCCL on Seq-Cifar-100 (CIL setting). Across buffer sizes, the additional time is a good trade-off, justified by significant performance gains. As in the discussion of [chen2025adaptiveretentioncorrection], other SOTA CL methods such as L2P [wang2022learningpromptcontinuallearning] and DualPrompt [wang2022dualpromptcomplementarypromptingrehearsalfree] increase training time by $37\%$ and $32\%$, respectively, for performance gains of only $2.5\%$ and $1.4\%$.
4.3.3 Performance with reduced batch size
To evaluate SAMix under smaller batches, we run TA-NCCL on Seq-Cifar-10 (Class-IL setting) with a batch size of 256 with different buffer sizes. As shown in Table 4, even without memory, SAMix with half the batch size outperforms the SOTA method CCLIS [li_cl_important_sampling_2024], which uses a batch size of 512 and memory size of 200. This highlights the strong adaptability of our method under constrained settings.
Table 3: Extra time (overhead, $\%$) and performance gain ($\%$) with SAMix on Seq-Cifar-100 (CIL setting)
Buffer 0 200 500
Overhead Gain Overhead Gain Overhead Gain
FC-NCCL 14.1 7.6 16.5 7.9 19.8 7.6
TA-NCCL 22.3 10.9 21.7 10.2 28.1 14.1
Table 4: SAMix results with varying batch sizes on Seq-Cifar-100 (CIL setting). Bold indicates comparison purposes
   Buffer    Method    Batch size    AA
   0    TA-NCCL    512    44.41
   TA-NCCL    256    42.77
   200    CCLIS    512    42.39
   TA-NCCL    512    46.95
   TA-NCCL    256    45.21